OpenStudy (anonymous):

Integrate from 2pi to infinitive. Sin tetha dtetha

OpenStudy (anonymous):

\[\int\limits\limits_{2 \pi}^{\infty}\sin \theta d \theta=-\cos \theta|_{2 \pi}^{\infty}\]Now we have a problem. The cosine curve oscillates indefinitely as we go to infinity. We cannot integrate this curve within these limits.